Solve the following system of equations.
x+9y=7 (1)
x=8-9y (2)
I'm either looking for an ordered pair or no solution at all.
I just cannot figure out what the (1) and (2) is doing there.

Answer by oberobic(2304)

x+9y = 7
x = 8-9y
.
x + 9y = 7
x + 9y = 8
.
These are parallel lines because there is no way that 0 = -1.
.
With parallel lines, there is no intersection.